Find the moment about the x-axis of a wire of constant density that lies along The curve

𝑦 = √π‘₯ from π‘₯ = 0 π‘‘π‘œ π‘₯ = 2


1
Expert's answer
2022-05-19T15:10:51-0400
"M_x=ρ \int_a^b \frac 1 2 ([f(x)]^2-[g(x)]^2) dx"

where


"a=0, b=2""f(x)=\sqrt {x}, \space g(x)=0""ρ - constant \space density"

then

"M_x=ρ \int_0^2 \frac 1 2 x dx=\frac ρ 2 \int_0^2 x dx=\frac ρ 2 \frac {x^2} 2 \bigg|^2_0=\frac ρ 2 \frac 4 2=ρ"

Answer: the moment about the x-axis is equal to the constant density: "M_x=\u03c1"